Let X and Y be independent random variable with means 3 ,5 and variances 16,25 respectively. Find the mean and standard deviations of 5X-3Y?


1
Expert's answer
2021-07-08T08:10:18-0400

Solution:

"\mu_X=3,\mu_Y=5
\\\sigma_X^2=16, \sigma_Y^2=25"

"\Rightarrow \sigma_X=4, \sigma_Y=5"

For "5X-3Y":

Mean"=E[5X-3Y]=5E[X]-3E[Y]=5(3)-3(5)=0"

Variance"=Var[5X-3Y]=25Var[X]+9Var[Y]-2(5)(3)Cov(X,Y)"

"Cov(X,Y)=0" for independent random variable.

Thus, "Var[5X-3Y]=25(4)+9(5)-0=100-45=55"

Then, standard deviation"=\sqrt{55}"